# BLAZE: Cross-Language and Cross-Project Bug Localization via Dynamic Chunking and Hard Example Learning

### v1

@@ Abstract
Software bugs require developers to exert significant effort to identify and resolve them, often consuming about one-third of their time. Bug localization, the process of pinpointing the exact source code files that need modification, is crucial in reducing this effort. Existing bug localization tools, typically reliant on deep learning techniques, face limitations in cross-project applicability and effectiveness in multi-language environments.
Recent advancements with Large Language Models (LLMs) offer detailed representations for bug localization. However, they encounter challenges with limited context windows and mapping accuracy. To address these issues, we propose BLAZE, an approach that employs dynamic chunking and hard example learning. First, BLAZE dynamically segments source code to minimize continuity loss. Then, BLAZE fine-tunes a GPT-based model using challenging bug cases, in order to enhance cross-project and cross-language bug localization. To support the capability of BLAZE, we create the BeetleBox dataset, which comprises 26,321 bugs from 29 large and thriving open-source projects across five different programming languages (Java, C+++++ +, Python, Go, and JavaScript). Our evaluations of BLAZE on three benchmark datasets—BeetleBox, SWE-Bench, and Ye et al.—demonstrate substantial improvements compared to six state-of-the-art baselines. Specifically, BLAZE achieves up to an increase of 120% in Top 1 accuracy, 144% in Mean Average Precision (MAP), and 100% in Mean Reciprocal Rank (MRR). An extensive ablation study confirms the contributions of our pipeline components to the overall performance enhancement.
@@ I Introduction
Recent studies have highlighted that Large Language Models (LLMs) provide detailed and context-rich representations for code retrieval tasks and bug diagnosis [17, 18]. This improved representation can enhance the ability of bug localization tools to accurately link bug reports with the corresponding source code files by integrating context-sensitive semantic understanding [19, 20]. However, such methods still encounter two major challenges. Firstly (Challenge 1: Limited Context Window), the typical context window of LLMs, which ranges from 512 to 2,048 tokens, is insufficient when compared to the average token length of 2,536 for a single file in common datasets [21].
@@ 1 Function dynamic_chunk(cost_map, total_lines, window_size)
Learning from hard examples. We fine-tune a Large Language Model (LLM), called CodeSage [31]. CodeSage is a GPT-based multi-modal embedding model. We adopt CodeSage to align bug reports and source code files within the same embedding space, drawing on its proven effectiveness in zero-shot code search tasks [31]. To optimize CodeSage for bug localization, we implement contrastive learning, which helps to distinguish between similar and dissimilar data pairs, i.e., this technique adjusts the model to minimize the distance between embeddings of bug reports and their corresponding buggy files (positive) while maximizing the separation from non-buggy files (negative). This distinction is guided by the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss, defined as:
@@ III.A Research Questions
A plethora of bug localization tools have been proposed (e.g.,  [12, 42, 43, 16]), with cross-project tools (e.g., [15, 16]) being particularly effective. However, these tools still require project-specific additional training. Our proposed approach, BLAZE, is designed to function without this requirement. This RQ evaluates how well BLAZE performs compared to state-of-the-art cross-project bug localization tools.
@@ III.B Dataset Curation of BeetleBox
  • Selection of repositories. We use the GitHub API search feature to collect repository names. We search for repositories written in the top five programming languages, Java, Python, C+++++ +, JavaScript, and Go, as ranked in the StackOverflow Survey. [^2] https://survey.stackoverflow.co/2022/#technology-most-popular-technologies We select the top 10 repositories for each language, sorting them in descending order by their star ratings and most recent updates. This step guarantees that the repositories are actively contributed to and popular within the open-source community.
@@
  • Filtration of issues. An issue may have several PRs, but typically only one is merged. To minimize errors in our ground truth, we filter out issues with the ‘duplicate’ tag and verify using GitHub API that each issue is linked to a single, merged PR. We initially started with 60 repositories. However, in certain projects, we found no issues meeting all the criteria. In the end, we are left with 29 repositories across five programming languages.
@@ IV Results
Approach. We train our model using BeetleBox. We conduct the training on 13K bugs from five languages, covering 13 projects. Our evaluation dataset contains 31,745 bugs from five languages covering 29 real-world software projects. Following prior studies [24, 48], BeetleBox uses a completely different set of projects for training and testing splits. As previously presented in Table I, the SWE-Bench and Ye et al. datasets are used only for testing.
@@ V Ablation Study
This section presents an ablation study to assess the influence of individual components in the BLAZE pipeline on its overall performance. Key strategies contributing to the effectiveness of BLAZE include (1) a combined pipeline of lexical and deep retrievers, (2) learning from hard examples, and (3) dynamic chunking. Next, we detail the results of the ablation study to evaluate the impact of each component.
@@ VI Threats to Validity
External validity. A threat to the external validity of our study lies in the generalizability of the results. We assess cross-project capabilities using separate projects for training and testing. Although the choice of test projects may affect performance, this approach aligns with previous studies that evaluate zero-shot and cross-project capabilities [48, 24]. Additionally, our evaluation of 31,745 real-world bugs ensures that our findings hold, even with different project selections, due to the large scale of our dataset. To address potential selection bias in the BeetleBox dataset, we collected project repositories based on star ratings and recent updates, ensuring that the repositories are actively contributed to and popular within the open-source community. This approach aims to include high-quality, widely used projects across the selected top five programming languages.
@@ VII Related Works
Other work proposed deep learning-based approaches to localizing bugs [53, 43, 5, 60, 11]. For example, FLIM [53] used CodeBERT embedding for calculating the similarity between bug reports and source code files at the function level. DNNLoc [61] combined a deep neural network with the VSM model to increase effectiveness across different types of metrics such as textual similarity, class name similarity, bug-fixing recency, and frequency. NP-CNN [43] employed a Convolutional Neural Network (CNN) to leverage both lexical and semantical properties. BL-GAN [12] proposed a generative adversarial network-based solution for bug localization. Several studies [42, 9, 8, 10] incorporated program structure information in the model to generate more useful program representation. For example, Liang et al. [10] and Li et al. [8] used abstract syntax trees, data flow graphs, and program flow graphs to generate a better representation of source code. While BLAZE employs an embedding model similar to the techniques discussed previously, it differs by its ability to be applied to new, unseen projects without requiring additional training.

### v2

@@ Abstract
Software bugs require developers to expend significant effort to identify and resolve them, often consuming about one-third of their time. Bug localization, the process of pinpointing the exact source code files that need modification, is crucial in reducing this effort. Existing bug localization tools, typically reliant on deep learning techniques, face limitations in both cross-project applicability and multi-language environments.
Recent advancements with Large Language Models (LLMs) offer detailed representations for bug localization that may help to overcome such limitations. However, these models are known to encounter challenges with 1) limited context windows and 2) mapping accuracy. To address these challenges, we propose BLAZE, an approach that employs dynamic chunking and hard example learning. First, BLAZE dynamically segments source code to minimize continuity loss. Then, BLAZE fine-tunes a GPT-based model using complex bug reports in order to enhance cross-project and cross-language bug localization. To support the capability of BLAZE, we create the BeetleBox dataset, which comprises 26,321 bugs from 29 large and thriving open-source projects across five programming languages (Java, C+++++ +, Python, Go, and JavaScript). Our evaluation of BLAZE on three benchmark datasets—BeetleBox, SWE-Bench, and Ye et al.—demonstrates substantial improvements compared to six state-of-the-art baselines. Specifically, BLAZE achieves up to an increase of 120% in Top 1 accuracy, 144% in Mean Average Precision (MAP), and 100% in Mean Reciprocal Rank (MRR). Furthermore, an extensive ablation study confirms the contributions of our pipeline components to the overall performance enhancement.
@@ I Introduction
Recent studies have highlighted that Large Language Models (LLMs) provide detailed and context-rich representations for code retrieval tasks and bug diagnosis [17, 18]. This improved representation can enhance the ability of bug localization tools to accurately link bug reports with the corresponding source code files by integrating context-sensitive semantic understanding [19, 20]. However, such methods still encounter two major challenges. Firstly (Challenge 1: Limited Context Window), the typical context window of LLMs, which ranges from 512 to 2,048 tokens, is insufficient when compared to the average token length of 2,536 for a single source code file in common datasets [21].
@@
BLAZE localizes bugs at the file level, narrowing the search space for faster bug localization and seamless integration with version control systems [24, 25]. We chose file-level localization for BLAZE to avoid the complexity (e.g., tracking changes in finer granularity [26, 27], hierarchical dependency of changes [28]) and overhead associated with method or commit-level approaches. This balance enhances both accuracy and usability [29, 30]. Additionally, higher granularity bug localization often leads to more false positives, which negatively impacts developers’ experience [31].
@@ 1 Function dynamic_chunk(cost_map, total_lines, window_size)
Learning from hard examples. We fine-tune a Large Language Model (LLM), called CodeSage [39]. CodeSage is a GPT-based multi-modal embedding model. We adopt CodeSage to align bug reports and source code files within the same embedding space, drawing on its proven effectiveness in zero-shot code search tasks [39]. However, its effectiveness in bug localization is still limited, as shown in Figure 1(a). A crucial difference between CodeSage and BLAZE lies in their strategies for mining hard examples; BLAZE mines in the embedding space, while CodeSage targets hard examples from docstrings and code. Mining in the embedding space enables BLAZE to detect hard examples across different contexts, enhancing its ability to generalize across a variety of programming languages and projects. To optimize CodeSage for bug localization, we implement contrastive learning, which helps to distinguish between similar and dissimilar data pairs, i.e., this technique adjusts the model to minimize the distance between embeddings of bug reports and their corresponding buggy files (positive) while maximizing the separation from non-buggy files (negative). This distinction is guided by the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [40], defined as:
@@ III.A Research Questions
A plethora of bug localization tools have been proposed (e.g.,  [12, 51, 52, 16]), with cross-project tools (e.g., [15, 16]) being particularly effective. However, these tools still require project-specific additional training. Our proposed approach, BLAZE, is designed to function without such a requirement. This RQ evaluates how well BLAZE performs compared to state-of-the-art cross-project bug localization tools.
@@ III.B Dataset Curation of BeetleBox
  • Selection of repositories. We use the GitHub API search feature to collect repository names. We search for repositories written in the top five programming languages, Java, Python, C+++++ +, JavaScript, and Go, as ranked in the StackOverflow Survey. [^2] https://survey.stackoverflow.co/2022/#technology-most-popular-technologies We select the top 20 repositories for each language, sorting them in descending order by their star ratings and most recent updates. This step guarantees that the repositories are actively contributed to and popular within the open-source community.
@@
  • Filtration of issues. An issue may have several PRs, but typically only one is merged. To minimize errors in our ground truth, we filter out issues with the ‘duplicate’ tag and verify using GitHub API that each issue is linked to a single, merged PR. We initially started with 100 repositories. However, in certain projects, we found no issues meeting all the criteria. Furthermore, to avoid potential data leakage [57], we have to filter out 63 repositories. In the end, we are left with 29 repositories across five programming languages.
@@ IV Results
Approach. We train our model using BeetleBox on 13K bugs from five languages, covering 13 projects. Our evaluation dataset contains 31,745 bugs from five languages covering 29 real-world software projects. Following prior studies [32, 58], BeetleBox uses a completely different set of projects for training and testing splits. As previously presented in Table I, the SWE-Bench and Ye et al. datasets are used only for testing.
@@ V Ablation Study
This section presents an ablation study to assess the influence of individual components in the BLAZE pipeline on its overall performance. Key strategies contributing to the effectiveness of BLAZE include (1) a combined pipeline of lexical and deep retrievers, (2) learning from hard examples, and (3) dynamic chunking. For all the experiments in this ablation study, we use BeetleBox as it contains bugs from five different languages. Next, we detail the results of the ablation study to evaluate the impact of each component.
@@ VI Threats to Validity
External validity. A threat to the external validity of our study lies in the generalizability of the results. We assess cross-project capabilities using separate projects for training and testing. Although the choice of test projects may affect performance, this approach aligns with previous studies that evaluate zero-shot and cross-project capabilities [58, 32]. Additionally, our evaluation of 31,745 real-world bugs ensures that our findings hold, even with different project selections, due to the large scale of our dataset. To address potential selection bias in the BeetleBox dataset, we collected project repositories based on star ratings and recent updates, ensuring that the repositories are actively contributed to and popular within the open-source community. This approach aims to include high-quality, widely used projects across the selected top five programming languages. Additionally, due to the absence of a replication package, we replicated TRANP-CNN and CooBA based on their descriptions in the original study, which may lead to slight performance deviations. Nevertheless, after experimenting with various hyperparameters, we selected a set that achieves comparable performance to that reported in the original study.
@@ VII Related Works
Other work proposed deep learning-based approaches to localizing bugs [63, 52, 5, 70, 11, 71]. For example, FLIM [63] used CodeBERT embedding for calculating the similarity between bug reports and source code files at the function level. DNNLoc [72] combined a deep neural network with the VSM model to increase effectiveness across different types of metrics such as textual similarity, class name similarity, bug-fixing recency, and frequency. NP-CNN [52] employed a Convolutional Neural Network (CNN) to leverage both lexical and semantical properties. BL-GAN [12] proposed a generative adversarial network-based solution for bug localization. Several studies [51, 9, 8, 10] incorporated program structure information in the model to generate more useful program representation. For example, Liang et al. [10] and Li et al. [8] used abstract syntax trees, data flow graphs, and program flow graphs to generate a better representation of source code. While BLAZE employs an embedding model similar to the techniques discussed previously, it differs by its ability to be applied to new, unseen projects without requiring additional training.